### v2

# Danny Halawi4, Fred Zhang4, Philip E. Tetlock2,5

Abstract
[…]
footnotetext: Preprint.footnotetext: * Equal Contributionfootnotetext: Correspondence to benchmark@forecastbench.org.footnotetext: This work was funded by a grant from Open Philanthropy; the grant supports the maintenance and updating of this benchmark for three years. We thank Sam Glover and Molly Hickman for writing superforecaster prompts, Otto Kuusela for helpful discussions, and Harrison Durland and Victoria Schmidt for their research assistance. The views expressed in this paper do not necessarily reflect the views of the Federal Reserve Bank of Chicago or the Federal Reserve system.
1 Introduction
Forecasting the future is a challenging but necessary task (Armstrong, 2001, Tetlock and Gardner, 2015). Economic forecasts influence investment and hiring decisions (Christensen et al., 2018). And forecasts of the Covid-19 pandemic in early 2020 prompted local lockdowns to slow the spread of the virus Adam (2020). However, human forecasting is often expensive, time-consuming, applicable only in specific domains, and subject to concerns about human biases. Motivated by these limitations, recent work explores the use of machine learning (ML) models, especially large language models (LLMs), in automated forecasting (Hendrycks et al., 2021, Zou et al., 2022, Yan et al., 2024, Pratt et al., 2024, Phan et al., 2024, Fluri et al., 2024, Halawi et al., 2024).
To assess LLMs’ forecasting capabilities, prior work built static benchmarks of questions where the answer was known (resolved) after the knowledge cut-offs of the LLMs they test (Zou et al., 2022, Yan et al., 2024, Halawi et al., 2024). For example, “Will a nuclear weapon be detonated in 2023,” though resolved on Jan 1, 2024, is a valid out-of-sample question for testing a model with a known knowledge cut-off before the end of 2023.
This static evaluation methodology comes with three key drawbacks. First, as the knowledge cut-offs of frontier models are updated, static benchmarks become obsolete, necessitating further data-sourcing. This makes it difficult to continuously track and compare the top models in the field. Second, knowledge cut-offs are usually estimated using the time range of pre-training data. Such estimates may not be accurate, and post-training may inject further post-cutoff knowledge into the model, contaminating the test sets. Lastly, model developers face financial incentives to exaggerate their accuracy on benchmarks and claim that their models are state-of-the-art performers. While some fudging is easily identified, other subtle benchmark manipulation or overfitting is harder to catch, and some studies have shown that there is significant evidence for benchmark contamination and/or memorization in popular LLM models (Roberts et al., 2023, Elazar et al., 2024, Li et al., 2023).
[…]
Our initial benchmark consists of 1,000 standardized forecasting questions randomly sampled from a much larger real-time question bank. To establish baseline levels of performance, we record predictions from dozens of LLMs on the initial set, using methods like retrieval-augmentation to boost model performance (Lewis et al., 2020, Halawi et al., 2024). We independently elicit predictions from human forecasters of varying expertise, including a group of seasoned forecasters (superforecasters) (Tetlock and Gardner, 2015) who have consistently performed well in competitive forecasting tournaments. As questions resolve, we rate the submissions of registered models and the human comparison groups, updating our public leaderboard.
Our initial results indicate that state-of-the-art models, such as GPT-4o and Claude-3.5, perform only roughly as well as a simple median of forecasts from a survey of humans with no (or minimal) forecasting experience, even when the LLMs are augmented with news retrieval and prompt engineering, and when the models have access to forecasts from a human crowd (on market-based questions). The models also perform significantly worse than the median of expert forecasters (superforecasters) who answer the same questions.
These findings leave significant room for researchers to attempt to improve AI-based forecasting systems using innovative approaches, such as developing methods for continuously updating models with current events and enhancing LLMs to reason over extended time frames. To support progress in this area, we publish an auxiliary dataset of LLM and human forecasts, rationales, and accuracy for use in future LLM fine-tuning and testing.
1.1 Related work
Automated forecasting
ML systems that make accurate forecasts can help inform human decision-making (Hendrycks et al., 2021, Schoenegger et al., 2024a). Recent work studies the use of LLMs for automated forecasting Jin et al. (2021), Zou et al. (2022), Yan et al. (2024), Halawi et al. (2024), Pratt et al. (2024). These papers all build static benchmarks of resolved questions. A recent paper from Halawi et al. (2024) uses questions resolved between June 2023 and January 2024. Unfortunately, the latest LLMs, such as GPT-4o, have knowledge cut-offs past this point. This fact motivates our work to build a dynamically updating benchmark that can accurately evaluate frontier model performance.
In addition, Schoenegger and Park (2023) and Abolghasemi et al. (2023) compare the accuracy of GPT-4 and other LLMs to human forecasters. Schoenegger et al. (2024b) found that an ensemble of $12$ LLMs rivaled human forecasts in a forecasting tournament in 2023 with a small number of questions, but the results are statistically imprecise. Unlike our work, that tournament was run only once and is no longer updated. Also, our much larger question set allows us to make precise statistical claims about the performance of LLMs relative to each other and relative to human performers.
Finally, recent work focuses on the use of LLMs and transformer-based systems in statistical time-series forecasting (Nie et al., 2023, Gruver et al., 2023, Dooley et al., 2023, Rasul et al., 2023, Jin et al., 2024, Das et al., 2024, Goswami et al., 2024, Woo et al., 2024), but many of the most important forecasting questions do not have well-defined time series that can be used in standard statistical forecasting models (e.g., what is the probability that a nuclear weapon will be used offensively in the next ten years?). Our benchmark is more general, and evaluates automated forecasting performance across questions with and without underlying time series and historical baseline data.
Language model evaluation
Evaluating highly capable LLMs is a challenging task—with many models saturating key benchmarks soon after their release (Maslej et al., 2023, Owen, 2024) and with benchmarks potentially leaked to models’ training data (Magar and Schwartz, 2022, Sainz et al., 2023, Jacovi et al., 2023, Xu et al., 2024b, Balloccu et al., 2024, Jiang et al., 2024b, Zhang et al., 2024, Xu et al., 2024a). Our dynamic forecasting benchmark avoids both of these issues. First, automated forecasting is highly unsaturated: Halawi et al. (2024) showed that under simple zero-shot evaluation, frontier models such as GPT-4 are much less accurate than aggregates of human predictions. Second, our benchmark is dynamic. The final resolution of each question is only determined in the future and cannot be ‘leaked’ in any training data (by construction). This eliminates concerns of contamination.
2 Preliminaries
Forecasting
A forecasting question asks for a probabilistic prediction of a future event. A forecaster may assign probabilities to potential outcomes of the event. Forecasting platforms, including prediction markets, host a wide range of questions. Many questions are of public interest, such as “Who will be the next US president in the 2024 election?”
Metrics
For binary questions, we use the Brier score as the performance metric, defined as $(f-o)^{2}$, where $f\in[0,1]$ is the probabilistic forecast and $o\in\{0,1\}$ is the outcome. Lower Brier scores are better, and a score of $0.25$ is associated with the uninformed forecast of $0.5$. Brier scores are strictly proper, incentivizing truthful reporting from participants.
Models
We evaluate 17 LLMs on our initial benchmark: GPT-3.5-Turbo-Instruct (Brown et al., 2020), GPT-4 (OpenAI, 2023), GPT-4o, Llama-2-70B (Touvron et al., 2023), Llama-3-7B, Llama-3-70B, Mistral-7B, Mistral-8x7B (Jiang et al., 2024a), Mistral-8x22B, Mistral-Large, Qwen1.5-110B-Chat, Claude-2.1 (Anthropic, 2023), Claude-3-Haiku, Claude-3.5-Sonnet, Claude-3-Opus (Anthropic, 2024), Gemini 1.5 Flash and Gemini 1.5 Pro (Gemini Team, 2023).
We record model predictions under three different methods: zero-shot prompting, scratchpad prompting, and scratchpad prompting with retrieval augmentation. For the last setting, we use the retrieval infrastructure from Halawi et al. (2024) and provide relevant news articles to the models in-context to reason about. Additionally, only models with a context window larger than 8,000 tokens were evaluated under the retrieval setting due to the inclusion of news articles in some prompts.
3 Benchmark, leaderboard, and datasets
We maintain a question bank with a growing set of forecasting questions. Continuously adding questions to the question bank allows it to stay relevant and ensures that we have a large selection of unresolved questions to sample from.
Every night, our automated system updates the question bank with new questions and resolution values. We drop invalid, low-quality, and resolved questions, categorizing the remaining questions by topic. The process is fully automated as detailed in Section 3.1.
Every two weeks, we sample from the question bank and release question sets for teams interested in submitting their forecasts to the benchmark. We also survey a standard set of LLM-based models to allow for comparisons of performance over time. The forecasts that these groups submit are resolved continuously with daily updates to our leaderboard. We provide the resulting data on forecast questions and submissions to researchers. See Section 3.2 for details. Finally, we discuss the question resolution procedure in Section 3.3 and our leaderboard design in Section 3.4.
3.1 Updating the question bank
In an automated process that runs nightly at 0:00 UTC, questions are added to the question bank, resolution values are updated, and metadata generated, as shown in Figure 1. Details follow.
Daily, 0:00 UTCUpdate QuestionsQuestion FilesQuestion FilesUpdate ResolutionsResolution FilesResolution FilesUpdate MetadataMetadata FileMetadata FileQuestionBank
Figure 1: The automated system that pulls questions from forecasting platforms and datasets for question generation from data providers.
3.1.1 Updating questions and resolution values
We bring in questions from two broad types of sources: markets and datasets. [^1] Licensing information can be found in Section C.1. We select multiple, reliable sources from each type, ensuring the diversity of our benchmark. See Table 1 for details.
Markets
As one of our two main question sources, we rely on a handful of prediction markets and forecast aggregation sites that elicit human predictions on questions across a wide range of topics. [^2] Hereafter, for brevity, we refer to questions that come from both prediction markets and forecast aggregation sites as ”market” questions. We also refer to theses sources collectively as ”market” sources. When selecting questions from market sources to add to the question bank, we choose those with high levels of active human trading on these platforms (liquidity) as these questions tend to be of higher quality than those with low levels of human trading. [^3] See Table 7 for an example question pulled from a market source. We also store the latest market and resolution values for each question.
Datasets
As a second question source, we rely on 5 well-established and well-maintained datasets that track real-world events (e.g. ACLED (Raleigh et al., 2023), a geopolitical database that tracks worldwide conflict, including facts like the number of protests in Niger each month). With these dataset sources, we can generate questions using a fixed question template (e.g., “Will the number of protests in Niger increase by at least 10% over this month’s value by the resolution date?"). [^4] See Table 8 for an example question pulled from a dataset source.
Question Bank
Table 1 lists the sources of market and dataset questions, along with the number of questions available in our question bank, whence we sample questions for our benchmark runs. In addition to the main questions (with sample size $N$), we also construct additional combination questions by choosing pairs of questions within each source. We describe combination questions in more detail in Section 3.2.
Table 1: Question bank composition, grouped by source type (market or dataset).
Source URL N $\binom{N}{2}$
RFI randforecastinginitiative.org 15 105
Manifold Markets manifold.markets 392 76,636
Metaculus metaculus.com 800 319,600
Polymarket polymarket.com 534 142,311
Market Total 1,741 538,652
ACLED acleddata.com 3,150 4,959,675
DBnomics db.nomics.world 52 1,326
FRED fred.stlouisfed.org 166 13,695
Wikipedia wikipedia.org 335 55,945
Yahoo! Finance finance.yahoo.com 504 126,756
Dataset Total 4,207 5,157,397
Question Bank Total 5,948 5,696,049
3.1.2 Updating forecast question metadata
After we automatically collect forecasting questions and resolution values in our standardized question bank, the data get processed in several ways. This generates more information about the questions and creates another sampling option for the creation of the final question set. Following Halawi et al. (2024), we use gpt-3.5-turbo-0125 to categorize questions by subject and to filter out low-quality questions. We report the number of questions by category and source in Table 14, where we display the breakdown of the standard questions from our question bank ($N$ from Table 1) across question categories by source, after removing low-quality questions.
3.2 Generating the question set
Every other Sunday at midnight UTC, we release a set of 1,000 forecast questions for LLMs through the process outlined in Figure 2. We sample an equal number of questions from each data source, to ensure representativeness. Within each source, we then uniformly sample questions across all question categories, aiming for an equal distribution from each category within each source. This ensures that models cannot be overfit to a specific type of question or topic. Limiting the number of questions generated to 1,000 also ensures that costs for testing LLM systems on the benchmark are capped for development teams with fewer resources.
For questions derived from dataset sources, the distribution of resolution dates is as follows, as represented by days past the forecast due date: [7, 30, 90, 180, 365, 1,095, 1,825, 3,650]. In other words, the distribution of resolution dates spans from one week into the future to ten years into the future. For questions derived from market sources, we ask only for the probability that each question will resolve positively (will the event underlying the question occur, or not occur). Prior to a given market question’s resolution, performance is evaluated by calculating the squared distance between the forecasted value and the crowd forecast on the platform from the previous day. In the main public leaderboard we report performance using all questions—including unresolved questions—but we also separately report performance on resolved questions for interested readers.
Every other Sunday,0:00 UTCSample QuestionsQuestionBankLLM Question Set   (1,000 Questions)Public Question Set(200 Questions)
Figure 2: We sample from the question bank every two weeks to produce question sets.
For every data source, there are two types of questions, each comprising half of the question set. The first type is a standard forecasting question with a binary outcome, e.g. “Will inflation (core CPI) be above 3% next month?” We construct the second type, combination questions, by pairing two standard questions. For combination questions, we ask for forecasts on all Boolean combinations of the two questions (i.e. $P(Q1\cap Q2)$, $P(Q1\cap\neg Q2)$, …). Considering the extensive number of existing standard questions and potential combinations in our question bank (as we could potentially combine 3, 4, or more standard questions together), we effectively have access to millions of possible forecasting questions to sample from. We show the number of two-question combination questions in the question bank as it stands at time of writing in the right-hand column of Table 1.
Combination questions require forecasters to consider the covariance structure of different events, some of which are more independent than others. For instance, the best forecasts of whether the S&P500 (a key U.S. stock market index) will reach an all-time high and whether Spain will win the next Men’s World Cup are likely independent. However, the best forecast of whether the S&P500 will reach an all-time high and whether the U.S. will enter an economic recession must account for the likely strong correlation between these events.
To estimate baseline human and expert performance on these questions, we run surveys asking the general public and expert forecasters to predict on our question sets (see Section 4). The human question is comprised of 200 forecast questions sampled directly from the LLM set of 1,000 questions. [^5] We can evaluate humans and models on more than 200 questions because each data question asks about the value of a datapoint at several different time horizons; and we consider accuracy of the model or human forecaster at each time horizon to be a different forecast. We produce the question sets 10 days before the forecast due date to allow for time to create and run a human survey. LLMs receive their question set 24 hours before the due date, even though it was generated at the same time as the human question set. This gives us the ability to elicit forecasts, obtaining comparable forecast sets on the due date from both LLMs and humans who faced the same information environment. A detailed timeline is visible in Figure 3.
[…]
Figure 3: Timeline from question set generation to forecast due date, $D$.
3.3 Resolution
We resolve forecast sets nightly by gathering the latest information about which events have or have not occurred. All questions are ultimately resolved to the final true value.
Markets
Ground truth is not available for market questions before the events are known to have occurred or not occurred. Before events resolve, we use the platform prediction (an aggregate of human forecasts reported on the platform) from the preceding day as a proxy resolution value for unresolved questions. These values are updated nightly as described in Section 3.1.1. This allows us to score the complete set of forecast questions, incorporating all information available to people on a nightly basis. Once a market question is officially resolved, we score the forecast against the resolution value, creating a definitive score for the question across all forecast horizons.
Datasets
Datasets can be updated as new information becomes available. Thus, questions derived from datasets are continuously resolved to the value from the the latest available data.
Missing forecasts
We select 1,000 questions for the LLM question set to make the forecasting task impractical to complete manually within the 24 hour window after the question set is released. And we obligate all LLMs to forecast on all questions to ensure comparability of scores across models and human-based aggregates. So, when a model does not submit predictions on certain questions or time horizons, we consider that a model error and impute a naive value for the model to ensure comparability across models over time.
For market questions, as we only ask users to forecast the outcome of the question, missing forecasts on question outcomes are assigned the value of the crowd forecast on the forecast due date, day $D$ in Figure 3. Some may argue this is overly-generous, but we did not want teams to have a competitive advantage by simply scraping the market websites themselves.
For dataset questions, we impute the value $0.5$ (which represents a lack of information) to forecasts across all time horizons. Empirically, top models report valid forecasts on all questions and are not affected by this imputation procedure.
3.4 Leaderboards
We generate leaderboards, ranking models and humans by average overall score. The main leaderboard highlights the top forecasting submission across all questions and can by sorted by performance on the source type (market or dataset) and by score on resolved questions.
3.5 Datasets
[…]
General public forecast dataset
Every time we run the public survey outlined in Section 4, we provide multiple independent forecasts and rationales on every one of the 200 forecast questions and report the accuracy of the median public forecast. See Section B.2.1 for details.
Superforecaster forecast dataset
We provided a random subset of the benchmark questions to 39 expert forecasters with a strong track record of accurate performance on a diverse set of geopolitical questions. In addition to forecasts and rationale, the superforecasters provide pertinent information about their forecasting process, like search terms used and useful URLs consulted. See Section B.2.2.
LLM forecast dataset
Similar to the general public dataset, we prompt LLMs to produce forecasts on each of 1,000 forecast questions in the LLM question set. Their rationales are also included in the dataset whenever provided. See Section B.2.3 for details.
Question & resolutions dataset
In creating the benchmark, we have automated question creation and resolution from all of the sources outlined in Table 1. We provide these as a dataset that can be combined with the forecast datasets mentioned above. See Section B.1 for details on these datasets.
4 Human forecaster baseline
4.1 Public
We recruited 500 human forecasters via Prolific and Facebook advertising to participate as representatives of the general public. These human subjects completed a brief introductory survey to gather demographic information and evaluate performance on a few forecasting and comprehension tasks. We then drew 200 questions from the 1,000 question set provided to language models; human subjects completed a one-hour survey containing 20 random questions from this 200-item subset where they provided their forecasts and rationales for each question.
The number of responses per question in the 200-item subset varied to ensure representativeness across subject matters and sources; at least 40 responses were gathered per question, averaging 49 responses per question.
4.2 Superforecasters
We additionally recruited 39 expert forecasters with a strong track record of accurate performance on a diverse set of geopolitical questions (“superforecasters”) to participate in a 9-day forecasting experiment in which participants were prompted to give their individual forecasts for 20 random questions from the same 200-item subset of forecasting questions we provided to the general public. Roughly halfway through the 9-day experiment, participants were moved into a group forecasting stage in which we allowed these expert forecaster participants to see one another’s forecasts and rationales and to communicate with one another about each question. They were also given the opportunity to forecast on questions beyond the 20 questions assigned to them individually.
Because of the more limited quantity of the superforecasters, questions generally had fewer responses than in the public survey; a minimum of 3 forecasts were recorded for each question, with an average of 8 responses per question.
See Figure 4 for an example question from the human surveys. All data from human surveys are anonymized per IRB requirements.
Results
We evaluate human performance on the forecasting tasks by comparing predictions to the ground truth on resolved questions and comparing to community aggregates on unresolved questions. We find that the median public survey participant had an overall Brier score of 0.107 on the 200-item subset we drew from the LLM Question Set with 1,000 questions. The median superforecaster participant had an overall Brier score of 0.093 on this same subset of forecasting questions. See Section 5 and Table 2 for a comparison between human performance and LLM performance, as well as Appendix H for a top 50 leaderboard of human and LLM performance.
5 LLM baseline
In this section, we evaluate the forecasting capabilities of language models and report on the methodology and results.
5.1 Methodology
We evaluate a suite of instruction-following chat models without any additional fine-tuning (see Section 2 for details on the models). In each baseline, we prompt the model to generate a probabilistic forecast that the question will resolve to “Yes."
Baselines
We implement six baselines: (1) zero-shot prompting, (2) prompting with scratchpad instructions, (3) zero-shot prompting with crowd forecasts, (4) scratchpad prompting with crowd forecasts, (5) prompting with scratchpad instructions and retrieved news articles, and (6) aggregating predictions from multiple LLMs. Each baseline is described in more detail below.
  • 1
    Our first baseline, prompts the model zero-shot to generate a forecast directly without generating other content, such as intermediate thinking (Figure 6). By prompting the model to output its forecast directly, we assess raw forecasting capability without sensitivity to prompting strategies.
  • 2
    Our second baseline, prompts the model with scratchpad instructions (Nye et al., 2021) that outline a procedure the model should use to reason about the question (Figure 7). Our scratchpad prompt comes from (Halawi et al., 2024), which formed its prompts through a combination of analyzing the Brier score as prompt changes were made, and by adding language to fix common errors the LLMs would make, e.g., asking them to rephrase the question for understanding.
  • 3
    The question sets we provide to LLMs contain what we term freeze values for market questions, which are just the crowd forecast on the market the day the question set was created, as decsribed in Section 3.2. Our third baseline is the zero-shot prompt with freeze values.
  • 4
    Our fourth baseline is the scratchpad prompt with freeze values.
  • 5
    Since LLMs’ knowledge are not continuously updated, it is important to provide LLMs with up-to-date information relevant to the question (Zou et al., 2022). Our fifth baseline, scratchpad with information retrieval, uses the same scratchpad prompts that also contain retrieved news articles.
    The retrieval system is the same as described in Halawi et al. (2024): an LLM generates search queries for a news API, filters articles for relevancy, and summarizes the articles.
  • 6
    The aggregated prediction of forecasters (“crowd" prediction) is a strong benchmark. For instance, Metaculus (2023) shows that an ensemble of all forecasters consistently outperforms using just the top 5, 10, …, 30 best forecasters (based on past scores). In our final baseline, we aggregate the predictions generated by LLMs into an LLM “crowd" prediction.
    To produce the LLM crowd forecast, we query GPT-4o, Claude-3.5-Sonnet, and Gemini-1.5-Pro. We use three prompts crafted by superforecasters who were asked to write prompts that would help an LLM produce accurate forecasts (Figure 8, Figure 9, Figure 10). This results in $3\times 3=9$ forecasts per question. We then show 3 LLM crowd baselines using the median, geometric mean, and geometric mean of log odds (Satopää et al., 2014). See details at Appendix E.
5.2 Results
Comparing Humans and LLMs
In Panel A of Table 2, we show that superforecasters achieve an overall mean Brier score of $0.093$, significantly outperforming both the general public (Brier $=0.107$, $p<0.01$) and the top LLM performer on the 200-item subset (Claude 3.5 Sonnet: Brier $=0.111$, $p\leq 0.01$). [^6] See statistical note in Appendix G. The top-performing language models all had access to the (human) crowd forecast on market questions. The top-performing model without access to the crowd forecast on market questions was significantly less accurate than models with access to the human forecast. The comparison between humans and LLMs relies on the 200 questions forecasted by humans, which is a random sub-sample of the 1,000 questions in the question set provided to LLMs. [^7] Accuracy measures are based on more than 200 forecasts because human and LLM forecasters submitted multiple forecasts on each dataset question, one for each time horizon. The results presented here include forecasts over the 7-day and 30-day time horizons.
Comparing LLMs
Panel B excludes humans and evaluates LLMs on the entire question set ($N=1,000$ questions). Here we see a similar ranking of models, with GPT-4-Turbo now slightly outperforming Claude-3.5 Sonnet. Both of those models outperform GPT-4o and Gemini-1.5 Pro. As in Panel A, most of the top-performing models use the scratchpad prompt (Figure 7) and use as inputs the human crowd forecasts for market questions. Access to recent topical news related to the questions did not improve performance.
Table 2: Leaderboard with Top 10 Models
Panel A: Includes human comparison group and evaluates the set of 200 questions received by humans
Brier Score $\downarrow$
Model Organization
Information
provided
Prompt
Dataset
(N=210)
Market
(N=77)
Overall
(N=287)
Confidence
Interval
Pairwise
p-value
comparing
to No. 1
Pct. more
accurate
than No. 1
Superforecaster median forecast ForecastBench 0.135 0.051 0.093 [0.073, 0.112] 0%
Public median forecast ForecastBench 0.167 0.048 0.107 [0.09, 0.125] 0.009 30%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.155 0.067 0.111 [0.09, 0.132] 0.01 33%
GPT-4 OpenAI Freeze values Zero shot 0.169 0.056 0.113 [0.093, 0.132] 0.003 37%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Zero shot 0.155 0.074 0.114 [0.09, 0.139] 0.008 36%
Claude-3-5-Sonnet-20240620 Anthropic News with freeze values Scratchpad 0.16 0.07 0.115 [0.095, 0.135] 0.004 31%
GPT-4o OpenAI Freeze values Scratchpad 0.173 0.059 0.116 [0.098, 0.133] 0.002 30%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Zero shot 0.177 0.055 0.116 [0.097, 0.135] <0.001 37%
GPT-4o OpenAI News with freeze values Scratchpad 0.179 0.059 0.119 [0.101, 0.137] <0.001 42%
Mistral-Large-Latest Mistral AI Freeze values Scratchpad 0.163 0.081 0.122 [0.107, 0.137] <0.001 29%
Always 0.5 ForecastBench 0.25 0.165 0.208 [0.198, 0.218] <0.001 17%
Panel B: Excludes human comparison group and evaluates the entire set of 1,000 questions
Brier Score $\downarrow$
Model Organization
Information
provided
Prompt
Dataset
(N=2,737)
Market
(N=912)
Overall
(N=3,649)
Confidence
Interval
Pairwise
p-value
comparing
to No. 1
Pct. more
accurate
than No. 1
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Scratchpad 0.175 0.051 0.113 [0.107, 0.118] 0%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.179 0.047 0.113 [0.108, 0.119] 0.442 53%
GPT-4o OpenAI Freeze values Scratchpad 0.201 0.039 0.12 [0.115, 0.125] <0.001 45%
Gemini-1.5-Pro Google Freeze values Scratchpad 0.168 0.075 0.121 [0.116, 0.127] <0.001 42%
GPT-4o OpenAI News with freeze values Scratchpad 0.196 0.05 0.123 [0.117, 0.128] <0.001 41%
Claude-3-Opus-20240229 Anthropic Freeze values Zero shot 0.192 0.061 0.127 [0.12, 0.133] <0.001 44%
Gemini-1.5-Pro Google News with freeze values Scratchpad 0.172 0.082 0.127 [0.121, 0.132] <0.001 41%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Zero shot 0.199 0.055 0.127 [0.12, 0.133] <0.001 47%
Qwen1.5-110B-Chat Qwen Freeze values Scratchpad 0.176 0.078 0.127 [0.122, 0.133] <0.001 32%
GPT-4-Turbo-2024-04-09 OpenAI No Scratchpad 0.175 0.083 0.129 [0.124, 0.134] <0.001 7%
Always 0.5 ForecastBench 0.25 0.16 0.205 [0.202, 0.208] <0.001 22%
Notes:
1. We only show the performance of one model from each model series. The full leaderboard is available at www.forecastbench.org. Online results are updated nightly, so may be slightly different than the static version presented here.
2. For resolved questions, predictions are compared against ground truth, while for unresolved questions, predictions are compared to community aggregates.
3. The overall score is calculated as the average of the mean dataset Brier score and the mean market Brier score.
4. Pct. more accurate than No. 1: The percent of questions where this forecaster had a better overall score than the best forecaster (with rank 1).
5. Pairwise p-value comparing to No. 1 (bootstrapped): The p-value calculated by bootstrapping the differences in overall score between each model and the best forecaster (the group with rank 1) under the null hypothesis that there’s no difference.
As a particular failure mode, we find LLMs are specifically worse at combination questions. Although our human surveys did not explicitly ask for forecasts on combination questions, we bound human performance by assuming independence of the component of each combination question. This underestimates human accuracy because a human forecaster predicting the outcome of a combination question could account for independence between the permuted events. In Table 19, we present this comparison of human and LLM forecasts. We see that LLMs perform poorly on these combination questions, and including them in the benchmark widens the gap between human and LLM performance: superforecasters (Brier score of $0.068$) outperform the general public (Brier score of $0.083$) and the top LLM (Claude 3.5, Brier score of $0.119$) significantly. To benchmark the size of this gap in performance, the $0.051$ Brier score gap in performance between superforecasters and Claude 3.5 (the top-performing LLM) is significantly larger than the $0.031$ gap in performance between Claude 3.5 and Claude 2.1.
6 Discussion
[…]
From a practical point of view, we hope that our benchmark can shed light on how automated systems compare in accuracy to humans and when they might become the preferred source of predictions. Finally, from the perspective of LLM evaluation, forecasting requires models to integrate pre-training knowledge, engage in effective reasoning, and maintain calibration to avoid overconfidence. All of these are essential capabilities that are desired across various domains. We believe that performance on our benchmark can shed light on other capabilities of LLMs.
Appendix A Licensing and maintenance
Hosting
[…]
Datasets
Our datasets, distributed under the CC BY-SA 4.0 license, will be made available and updated regularly on www.forecastbench.org.
Codebase
The code underlying our automated system runs on Google Cloud Platform and is available at https://github.com/forecastingresearch/forecastbench under the MIT license.
Maintenance & long-term preservation
We ensure the long-term availability and maintenance of the benchmark as it is funded by Open Philanthropy until mid-2027. In case no further funding is provided beyond that point, we will upload our final datasets to GitHub and Hugging Face.
Appendix B Datasets
We intend our datasets to be used for training general LLMs, fine-tuning forecasting LLMs, and for any applicable research purposes. No restrictions are placed on who may use our datasets, nor to what end.
B.1 Question and resolution sets
Every question set will be published. Their resolutions will also be published such that there’s a complete training set when combined with the forecast sets outlined in Section B.2. The data dictionary for the question set is outlined in Table 3 and Table 4. The data dictionary for the resolution set is outlined in Table 5 and Table 6.
Data format
The question and resolution datasets are released as JSON (.json) files.
Ethical and responsible use
There are no restrictions on use of the question and resolution datasets.
Availability
Data collection
Our question and resolution datasets have been pulled, and are updated, from various, public-facing sources. From those sources where the terms of use/service prohibit the redistribution of their information (currently, Manifold Markets and Metaculus), we have obtained explicit permission to do so. Before we add new sources to our growing dataset, we will ensure the ability to distribute questions and resolutions publicly. Data sources in our question bank can be found in Table 13.
Table 3: Question set data dictionary.
Field Description Required Data Type
forecast_due
_date
Date in ISO format. e.g. "2024-07-21" string
question_set The name of the file that contains the question set. e.g. "2024-07-21-llm.json" string
questions A list of questions to forecast on, as defined in Table 4. array<object>
Table 4: Data dictionary for question entries in questions array from Table 3.
Field Description Required Data Type
id A unique identifier string given source. If instead of a string it’s an array of strings, then this is a combination question and combination_of will contain one question per id in the array of strings. string | array<string>
source Where the data comes from. string
question The question to forecast, presented as an f-string with placeholders {forecast_due_date} and {resolution_date} for dataset questions. string
resolution_
criteria
[…]
background Background information about the forecast question provided by the source, if available. Default: ‘N/A’ string
market_info_
open_
datetime
The datetime when the forecast question went on the market specified by source. Default: ‘N/A’ string
market_info
_close_
datetime
The datetime when the forecast question closes on the market specified by source. Default: ‘N/A’ string
market_info_
resolution_
criteria
The resolution criteria provided by the source, if available. Default: ‘N/A’ string
url The URL where the resolution value is found. string
freeze_datetime The datetime UTC when this question set was generated. string
freeze_datetime_
value
The latest value of the market or comparison value the day the question was frozen. string
freeze_datetime_
value
_explanation
Explanation of what the value specified in freeze_datetime_value represents. string
source_intro A prompt that presents the source of this question. string
combination_of An array of question objects, as defined by this data dictionary. Default: ‘N/A’ string | array<object>
resolution_dates The resolution dates for which forecasts should be provided for this forecast question. array<string>
Table 5: Resolution set data dictionary.
Field Description Required Data Type
forecast_due
_date
Date in ISO format. e.g. "2024-07-21" string
question_set The name of the file that contains the question set. e.g. "2024-07-21-llm.json" string
resolutions A list of resolutions to the forecast questions, as defined in Table 6. array<object>
Table 6: Data dictionary for resolution entries in questions array from Table 5.
Field Description Required Data Type
id A unique identifier string given source. A list if this was a combination question. string | array<string>
source Where the data comes from. string
direction If id has an array value, this is an array of the same length. Each entry $\in\{-1,1\}$. If the value is $1$, it means the question was asked in the normal direction. If the value is $-1$, it means the question was negated in the question. e.g. for a question asking for $P(\neg Q1\cap Q2)$, the value would be $[-1,1]$; all possible directions for this question would be: $[1,1]$, $[-1,1]$, $[1,-1]$, $[-1,-1]$. The value is null when id is a string. string
forecast_due
_date
The date the forecast was due in ISO 8601 format YYYY-MM-DD; day $D$ from Figure 3. string
resolution_date The date the value is associated with in ISO 8601 format YYYY-MM-DD. string
resolved_to The resolution value for the given date. number
resolved If true the question has been resolved. False otherwise. boolean
B.1.1 Examples
Table 7, Table 8, and Table 9 show concrete examples of the data dictionary detailed in Table 3.
Table 7: Prediction market example.
Field Entry
id 1558
source metaculus
combination_of N/A
question Will the Cavendish account for less than 50% of banana exports worldwide before 2035?
background Bananas are a well-liked import fruit all over the world, and the Cavendish cultivar has been crushing that market for sixty years. But its rise is literally founded upon the compost heap of the Gros Michel, another cultivar. The so-called ’Big Mike’ variety had been the leading export towards Europe and North America, but the Panama disease, a fungus belonging to the *Fusarium* clade, killed that. Luckily the Cavendish, grown in the same soil as the wilting Gros Michel, replaced it as *the* banana most of the western world connected with bananas. However, it appears another *Fusarium* rears its spores. Cavendish, with their genetic homogenity (they’re all clones) and sterile nature, aren’t resistant to it, and the fungus is ravaging more and more plantations. There are efforts under way to deal with *Fusarium*, but with various societies’ doubts and misgivings about GMOs, the cure may be viewed as a curse instead.
market_info_
resolution_
criteria
This question will resolve as **Yes** if the Cavendish banana accounts for less than 50% of worldwide annual banana exports in any year from 2018 to 2034 (inclusive).
market_info_
open_datetime
2018-11-13T08:00:00+00:00
market_info_
close_datetime
2034-12-31T23:00:00+00:00
resolution_
criteria
Resolves to the outcome of the question found at https://www.metaculus.com/questions/1558/cavendish-bananas-collapse-by-2035/.
url https://www.metaculus.com/questions/1558/cavendish-bananas-collapse-by-2035/
freeze_datetime_
value
0.48
freeze_datetime_
value_explanation
The community prediction.
freeze_datetime 2024-07-12T00:00:00+00:00
source_intro We would like you to predict the outcome of a prediction market. A prediction market, in this context, is the aggregate of predictions submitted by users on the website Manifold. You’re going to predict the probability that the market will resolve as ‘Yes’.
resolution_dates N/A
Table 8: Data source (DBnomics) example.
Field Entry
id meteofrance_TEMPERATURE_celsius.07005.D
source dbnomics
combination_of N/A
question What is the probability that the daily average temperature at the French weather station at Abbeville will be higher on resolution_date than on forecast_due_date?
background The history of Average temperature by day and by station for France - Degree Celsius - ABBEVILLE - Daily from Météo-France is available at https://db.nomics.world/meteofrance_TEMPERATURE_celsius.07005.D.
market_info_
resolution_
criteria N/A
market_info_
open_datetime N/A
market_info_
close_datetime N/A
url https://db.nomics.world/meteofrance_TEMPERATURE_celsius.07005.D
resolution_
criteria
Resolves to the value found at https://db.nomics.world/meteofrance_TEMPERATURE_celsius.07005.D once the data is published.
freeze_datetime_
value 17.95
freeze_datetime_
value_explanation
The daily average temperature at the French weather station at Abbeville.
freeze_datetime 2024-07-12T00:00:00+00:00
source_intro DBnomics collects data on topics such as population and living conditions, environment and energy, agriculture, finance, trade and others from publicly available resources, for example national and international statistical institutions, researchers and private companies. You’re going to predict how questions based on this data will resolve.
resolution_dates ["2024-07-28", "2024-08-20", "2024-10-19", "2025-01-17", "2025-07-21", "2027-07-21", "2029-07-20", "2034-07-19"]
Table 9: Combination (DBnomics) example.
Field Entry
id ["meteofrance_TEMPERATURE_celsius.07117.D", "meteofrance_TEMPERATURE_celsius.07240.D"]
source dbnomics
combination_of [An array containing dictionary entries of both questions.]
question We are presenting you with two probability questions. Please predict the probability that both will happen, that one will happen but not the other, and that neither will happen. In other words, for each resolution date please provide 4 predictions.
background N/A
market_info_
resolution_
criteria N/A
market_info_
open_datetime N/A
market_info_
close_datetime N/A
url N/A
resolution_
criteria N/A
freeze_datetime
value N/A
freeze_datetime
value_
explanation N/A
freeze_datetime 2024-07-12T00:00:00+00:00
human_prompt We are presenting you with two probability questions. Please predict the probability that both will happen, that one will happen but not the other, and that neither will happen. In other words, for each resolution date please provide 4 predictions.
resolution_dates ["2024-07-28", "2024-08-20", "2024-10-19", "2025-01-17", "2025-07-21", "2027-07-21", "2029-07-20", "2034-07-19"]
B.2 Forecast sets
[…]
Each forecast set contains the header information outlined in Table 10, with all forecasts in an array called forecasts. The forecast sets are described in the following subsections.
Data format
The question and resolution datasets are released as JSON (.json) files.
Ethical and responsible use
There are no restrictions on use of the forecast sets.
Availability
Table 10: Data dictionary of headers for forecast set.
Field Description Required Data Type
organization The organization name as it should be displayed on the leaderboard. string
model The model name as it should be displayed on the leader board. string
question_set The name of the question set file these forecasts are associated with. string
forecast_due
_date
The date the forecasts were due in ISO 8601 format YYYY-MM-DD. string
forecasts All forecasts for this question set. array<object>
B.2.1 General public forecast set
This forecast set consists of both forecasts made by individuals on the given question set and of the aggregation of those forecasts, as described in Appendix D. The dataset will be updated every time a survey is run with every forecast containing the information outlined in Table 10 and Table 11. Note that the aggregated forecast sets do not contain the user_id or reasoning fields.
Data collection
Data is collected from human forecasters via Qualtrics. We collect the reasoning behind forecasts and manually anonymize the data to ensure there is no personally identifiable information in the dataset before posting it online.
Table 11: Public forecast set data dictionary of entries in forecasts array from Table 10.
Field Description Required Data Type
id A unique identifier string given source. If instead of a string it’s a list of strings, then this is a combination question. string | array<string>
source Where the data comes from. string
forecast The forecast $\in[0,1]$. number
resolution_date The resolution date this forecast corresponds to. null for market questions. string |null
reasoning The rationale underlying the forecast. string
direction If id has an array value, this is an array of the same length. Each entry is an integer $\in\{-1,1\}$. If the value is $1$, it means the question was asked in the normal direction. If the value is $-1$, it means the question was negated in the question. e.g. for a question asking for $P(\neg Q1\cap Q2)$, the value would be $[-1,1]$. All possible values are: $[1,1]$, $[-1,1]$, $[1,-1]$, $[-1,-1]$, and null. array<number> |null
user_id A randomly generated string associated with the human respondent who submitted the forecast. This value contains no information which could identify said participant and was assigned to the dataset after personal identifiers had been removed. string
B.2.2 Superforecaster forecast set
In addition to the fields outlined in Table 10 and Table 11, the superforecaster dataset contains forecasts with the fields shown in Table 12. Likewise, this dataset will be updated every time a group of superforecasters is surveyed.
Data collection
39 superforecasters provided forecasts, rationales, and additional information about they way they forecast in our latest survey round. We will further manually check all of their responses to ensure the anonymity of the dataset.
Table 12: Superforecaster forecast set data dictionary of entries in forecasts array from Table 10 (additional fields to those in Table 11).
Field Description Required Data Type
searches An array of search terms used in researching the topic. array<string>
consulted_urls A list of useful urls array<string>
B.2.3 LLM forecast set
This dataset provides the same data (aside from user_id) as outlined in Table 10 and Table 11, only provided by language models. Each individual .json file was created by a model for the given question set.
Data collection
Beyond informing teams their forecasts will be made public, we will not check the rationales.
Appendix C Question Bank
C.1 Licenses
The licenses outlined in Table 13 apply to the datasets we have sourced for questions. We were granted express permission to use questions from Manifold Markets and Metaculus. Though not required by their license, we also met with a representative from ACLED who approved our use of their dataset for the benchmark and dataset distribution.
Table 13: Question sources and permissions
License License grants permission to use Express permission granted
RAND Forecasting Initiative Public Domain
Manifold Terms of Service
Metaculus Terms of Use
Polymarket Terms of Service
ACLED Terms of Use
DBnomics Open License
FRED Terms of Use
Wikipedia Terms of Use
Yahoo! Terms of Use
C.2 Categories
We generate metadata on all of the questions in our question bank, categorizing our questions, as described in Section 3.1.2.
It is important to have forecasting questions across a broad array of categories to test LLM capabilities. We are currently adding more questions from datasets with the goal of equilibrating these categories.
Table 14: Categories and question counts by source, dropping invalid questions.
RFI
Manifold
Metaculus
Polymarket
ACLED
DBnomics
FRED
Wikipedia
Yahoo!
Total
Arts & Recreation 0 57 10 16 0 0 0 0 0 83
Economics & Business 1 22 76 35 0 0 166 0 504 804
Environment & Energy 0 3 44 5 0 52 0 0 0 104
Healthcare & Biology 0 8 83 1 0 0 0 214 0 306
Politics & Governance 6 51 208 428 0 0 0 0 0 693
Science & Tech 5 101 200 21 0 0 0 1 0 328
Security & Defense 4 15 157 9 3,150 0 0 0 0 3,335
Sports 0 77 21 4 0 0 0 120 0 222
Other 0 0 1 0 0 0 0 0 0 1
Total 16 334 800 519 3150 52 166 335 504 5,876
Appendix D Human Survey Instructions and Screenshots
D.1 Instructions
Participants in the public survey were first prompted to read a consent form detailing the tasks involved in the study and potential risks to participants, alongside estimates of payment for completing each study (\$5 + bonus payments for forecasting performance in the introductory survey; \$10 for completing the primary survey). Then, participants were prompted to give their responses to 20 forecasting questions randomly selected from the 200-question subset of questions provided to the LLMs.
Participants were given a brief description of the task ahead before each question as follows:
You are going to be predicting the probability of the answer to the question below being "Yes" (or "resolving positively").
For tasks with questions generated from data providers, forecasters were prompted to provide a probability forecast at multiple resolution dates.
Participants in the superforecaster survey went through a similar initial experience before being moved into the group forecasting stage described in Section 4. Because of the expertise of this set of forecasters, superforecaster participants were also guaranteed a base payment of \$1,000 for participating in the individual and group stages of the experiment as well as bonus payments for individual-level accuracy.
D.2 Screenshots
The following screenshots (Figure 4, Figure 5 show a few of the questions presented to participants in the public study.
Figure 4: An example market-based question from the human survey.
Figure 5: An example question generated from a data provider, in this case DBnomics, from the public survey. Two of eight time horizons for which we elicited forecasts are included above.
Appendix E Details about LLM "Crowd" Baseline
E.1 Models
To construct a crowd baseline that includes diverse candidates, we evaluate models using the most recent forecasting dataset containing cross-domain questions with true resolutions from Halawi et al. (2024). We assess models from the following organizations: OpenAI, Mistral AI, Qwen, Google, Anthropic, and Meta. Using the same scratchpad prompting method from Halawi et al. (2024), we then select the top three models: GPT-4o, Gemini-1.5.Pro, Claude-3.5-Sonnet. See Table 15 for the results.
Table 15: Brier Scores from each LLM "crowd" candidate.
Model Scratchpad
GPT-4o 0.207 (0.026)
Llama-3-70b 0.232 (0.020)
Mistral-Large 0.233 (0.026)
Qwen-1.5-110b 0.222 (0.025)
Gemini-1.5-Pro 0.214 (0.025)
Claude-3.5-Sonnet 0.178 (0.025)
GPT-4-0613 0.222 (0.009)
GPT-4-1106-Preview 0.209 (0.012)
GPT-3.5-Turbo-1106 0.261 (0.010)
GPT-3.5-Turbo-Instruct 0.257 (0.009)
Claude-2 0.219 (0.014)
Claude-2.1 0.215 (0.014)
Gemini-Pro 0.230 (0.007)
Mistral-7B-Instruct 0.243 (0.008)
Mistral-8x7B-Instruct 0.238 (0.010)
Mixtral-8x7B-DPO 0.248 (0.010)
Yi-34B-Chat 0.241 (0.009)
Llama-2-7B 0.264 (0.011)
Llama-2-13B 0.268 (0.008)
Llama-2-70B 0.282 (0.011)
The highlights indicate the models we decide to use. Subscript numbers denote 2 standard errors. Above the divider line are the new models we evaluate, while below the divider line are the models evaluated by Halawi et al. (2024) on the same dataset using the same prompt.
E.2 Aggregation Methods
With the forecasts generated by the top 3 models selected in Section E.1, we then compare performance of 5 aggregation methods: Median, Trimmed Mean, Geometric Mean, and Geometric Mean of Log Odds (Satopää et al., 2014). See Table 16 for the results.
Table 16: Brier Scores from each Aggregation Method.
Aggregation Method Brier Score
Median 0.197 (0.024)
Trimmed Mean 0.197 (0.024)
Geometric Mean 0.194 (0.023)
Geometric Mean of Log Odds 0.194 (0.023)
The geometric mean and geometric mean of log odds are highlighted as they achieve the best performance. Subscript numbers denote 2 standard errors.
E.3 Other Details
LLM Parameters. We set the temperature as 0 and the max output token length as 2000.
Appendix F Prompts
In this section, we present the following prompts: zero-shot (Figure 6), scratchpad (Figure 7), and three prompts (Figure 8, Figure 9, Figure 10) written by superforecasters from the Forecasting Research Institute used to construct the LLM "crowd" baseline. Note that the scratchpad with information retrieval prompt is simply the scratchpad with an additional line "We have retrieved the following information for this question: {retrieved_info}" before the instructions begin. For combination questions, we slightly modify each of the above prompts by including the second question’s information (Figure 11).
{mdframed}
You are an expert superforecaster, familiar with the work of Tetlock and others. Make a prediction of the probability that the question will be resolved as true. You MUST give a probability estimate between 0 and 1 UNDER ALL CIRCUMSTANCES. If for some reason you can’t answer, pick the base rate, but return a number between 0 and 1.
Question: {question}
Question Background: {background}
Resolution Criteria: {resolution_criteria}
Question close date: {close_date}
Output your answer (a number between 0 and 1) with an asterisk at the beginning and end of the decimal. Do not output anything else. Answer:
{{ Insert answer here }}
Figure 6: Zero-shot Prompt from Halawi et al. (2024)
{mdframed}
Question:
{question}
Question Background:
{background}
Resolution Criteria:
{resolution_criteria}
Question close date: {close_date}
Instructions:
1. Given the above question, rephrase and expand it to help you do better answering. Maintain all information in the original question.
{{ Insert rephrased and expanded question.}}
2. Provide a few reasons why the answer might be no. Rate the strength of each reason.
{{ Insert your thoughts }}
3. Provide a few reasons why the answer might be yes. Rate the strength of each reason.
{{ Insert your thoughts }}
4. Aggregate your considerations. Think like a superforecaster (e.g. Nate Silver).
{{ Insert your aggregated considerations }}
5. Output an initial probability (prediction) given steps 1-4.
{{ Insert initial probability }}
6. Evaluate whether your calculated probability is excessively confident or not confident enough. Also, consider anything else that might affect the forecast that you did not before consider.
{{ Insert your thoughts }}
7. Output your answer (a number between 0 and 1) with an asterisk at the beginning and end of the decimal. (For example, if there are n resolution dates, you would output different *p* for each resolution date) Do not output anything else.
{{ Insert your answer }}
Figure 7: Scratchpad Prompt modified from Halawi et al. (2024)
{mdframed}
Question: {question}
Question Background: {background}
Resolution Criteria: {resolution_criteria}
Question close date: {close_date}
We have retrieved the following information for this question: {retrieved_info}
Instructions:
1. Given the above question, rephrase and expand it to help you do better answering. Maintain all information in the original question. {{ Insert rephrased and expanded question.}}
2. Let’s start by coming up with a base-rate that could be helpful for forecasting this question. Come up with the best reference-class you can for this sort of event, and give a general base-rate that doesn’t take into account factors unique to this question.
For instance, if the question were about the probability of a new technology being widely adopted within five years, you might look at historical data on the adoption rates of similar technologies as a reference class. Come up with a base-rate that could be relevant for this question.
The base-rate must be formatted as a clear probability (or number, in cases where you believe that to be more useful than a probability). For instance, imagine you are forecasting the probability that an incumbent president will be re-elected in an upcoming election in a hypothetical country. The past data shows that the incumbent has been elected 60% of the time.
Here, you would write ‘The reference class I have chosen is the incumbent being elected. My base-rate is that the probability of the incumbent being re-elected is 0.6.’ Give a justification for the base-rate, as well as a clear number.
Importantly, the base-rate should be as specific as it’s possible to be without losing confidence that the number is correct. For instance, if you were forecasting on the probability of a hypothetical democratic country going to war in the next year, you should ideally produce a base-rate for a democratic country going to war in a given year, rather than simply thinking about a given country going to war.
{{ Insert your base rate }}
3. Now, let’s think about factors specific to this question that may give us a good reason to deviate from the base-rate. Please give some reasons that the probability of this question resolving positively may be higher than the base rate. Please note specifically how they affect your forecast in terms of percentage point change. {{ Insert your thoughts }}
4. Now, let’s think about reasons that the probability of this question resolving positively may be lower than the base rate. Please note specifically how they affect your forecast in terms of percentage point change. {{ Insert your thoughts }}
5. Consider any other factors that may affect the probability of this question resolving positively or negatively, that you have not already discussed in the previous two steps. {{ Insert your thoughts }}
6. Aggregate your considerations. Think like a superforecaster (e.g. Nate Silver). Give a ranking to each consideration based on how much you believe it ought to affect your forecast. {{ Insert your aggregated considerations }}
7. Are there any ways in which the question could resolve positively or negatively that you haven’t considered yet, or that require some outside-the-box thinking? For example, if the question was ‘Will Microsoft have a market capitalization of over \$5tn by 2030’, you might consider questions like:
How likely is it that Microsoft no longer exists in 2030? How likely is it that inflation erodes that value of the dollar as such that \$5n is worth significantly less than it is today? How likely is it that there is a merger between Microsoft and another large company? How likely is it that Microsoft is broken up, as it is perceived to have monopoly power?
Here, we’re thinking about things that are probably quite unlikely to happen, but should still be integrated into your forecast. Write up some possibilities and consider how they should be integrated into your final forecast. {{Insert your thoughts and considerations about how this should affect your forecast}}
8. Output an initial probability (prediction) given steps 1-7. {{ Insert initial probability. }}
9. Okay, now let’s think about some other ways to consider how to forecast on this question. What would you say are the odds that if you could fast-forward and find out whether that statement is true or false, you would find out it’s true? You must give an odds ratio. This odds ratio probably shouldn’t be purely on the basis of the considerations in the previous steps, but you should think again about what you would expect to see if you could fast-forward into the future. If it helps, imagine that you’re taking a bet. {{ Insert your odds ratio. }}
10. Given your rephrased statement from step 1, think of 2-3 statements that if you conditioned on their being TRUE, you would think it more or less likely that your statement would be TRUE as well. These statements must not DETERMINE OR BE LOGICALLY EQUIVALENT to the original statement. Be creative! {{ Insert 2 to 3 related statements. }}
11. For each of your related statements, give new odds of the original statement conditional on the related statement being TRUE. {{ For each related statement, insert new odds for the original statement. }}
12. Now consider each of your odds from the previous steps(steps 9 - 11), and come up with your all-things-considered odds ratio for the original statement. {{ Insert final odds for the original statement. }}
13. Now, convert that odds ratio to a probability between 0 and 1. {{Insert a probability}}
14. Now, consider the probability that you came up with in step 8, as well as the probability that you came up with in step 13. Which of these probabilities do you lean towards? How do you weigh them against one another? Write up your thoughts on which probability is more likely to be “correct”, and then decide on a FINAL probability that will be used as your forecast. {{Insert your thoughts AND a final probability}}
15. Output your answer (a number between 0 and 1) with an asterisk at the beginning and end of the decimal. {{ Insert your answer }}
Figure 8: Superforecaster Prompt 1
{mdframed}
Question: {question}
Question Background: {background}
Resolution Criteria: {resolution_criteria}
Here’s some related information from the news that I’ve collected for you: {retrieved_info}
Question close date: {close_date}
Instructions:
1. Rephrase the question as a statement about the future, e.g. you would rephrase “Will Biden be the U.S. president on January 1 2025?” as “Biden is the U.S. president on January 1 2025.” {{ Insert question rephrased as a statement. }}
2. What would you say are the odds that if you could fast-forward and find out whether that statement is true or false, you would find out it’s true? You must give an odds ratio. If it helps, imagine that you’re taking a bet. {{ Insert your odds ratio. }}
3. Given your rephrased statement, think of 2-3 statements that if you conditioned on their being TRUE, you would think it more or less likely that your statement would be TRUE as well. These statements must not DETERMINE OR BE LOGICALLY EQUIVALENT to the original statement. Be creative! {{ Insert 2 to 3 related statements. }}
4. For each of your related statements, give new odds of the original statement conditional on the related statement being TRUE.insert new odds for the original statement. }}
5. Now consider each of your odds from the previous steps and come up with your all-things-considered odds ratio for the original statement. Output your answer (a number between 0 and 1) with an asterisk at the beginning and end of the decimal. {{ Insert final odds for the original statement. }}
Figure 9: Superforecaster Prompt 2
{mdframed}
Question: {question}
Question Background: {background}
Resolution Criteria: {resolution_criteria}
Relevant information we retrieved from news articles: {retrieved_info}
Question close date: {close_date}
Instructions:
1. Given the above question, rephrase and expand it to help you do better answering. Maintain all information in the original question. {{ Insert rephrased and expanded question.}}
2. Provide a few reasons why the answer might be no. Rate the strength of each reason. For now, ignore the evidence, ideas, and perspectives contained in the attached news articles. {{ Insert your thoughts }}
3. Provide a few reasons why the answer might be yes. Rate the strength of each reason. For now, ignore the evidence, ideas, and perspectives contained in the attached news articles. {{ Insert your thoughts }}
4. Aggregate the considerations you developed in the previous steps. Think like a superforecaster (e.g. Nate Silver). {{ Insert your aggregated considerations }}
5. Output an initial probability (prediction) given steps 1-4. {{ Insert initial probability. }}
6. Now, consider the perspectives, ideas, and evidence that was provided in the retrieved news articles. How should these affect your judgment of the probability of the question resolving positively? List all reasons why these news articles might increase the probability of the question resolving positively. {{Insert your thoughts}}
7. Now, let’s focus on how the ideas, perspectives, and evidence provided in the news articles might decrease the probability of the question resolving positively. {{Insert your thoughts}}
8. Given what you’ve thought about in the previous two steps, update your probability from the initial probability you gave in step 5. {{Insert updated probability}}
9. Evaluate whether your calculated probability is excessively confident or not confident enough. Also, consider anything else that might affect the forecast that you did not before consider. {{ Insert your thoughts }}
10. Output your answer (a number between 0 and 1) with an asterisk at the beginning and end of the decimal. Do not output anything else. {{ Insert your answer }}
Figure 10: Superforecaster Prompt 3
{mdframed}
Question 1: {question_1}
Question 2: {question_2}
Question 1 Background: {background_1}
Question 2 Background: {background_2}
Question 1 Resolution Criteria: {resolution_criteria_1}
Question 2 Resolution Criteria: {resolution_criteria_2}
Question 1 Current value on {freeze_datetime_1}: {value_at_freeze_datetime_1}
Question 1 Value Explanation: {value_at_freeze_datetime_explanation_1}
Question 2 Current value on {freeze_datetime_1}: {value_at_freeze_datetime_2}
Question 2 Value Explanation: {value_at_freeze_datetime_explanation_2}
Here’s some related information from the news that I’ve collected for Question 1: {retrieved_info_1}
Here’s some related information from the news that I’ve collected for Question 2: {retrieved_info_2}
Question resolution date: {list_of_resolution_dates}
Figure 11: Combination Prompt that includes information about both non-market questions. The instructions are truncated and can be supplemented with any of the prompts shown above.
Appendix G Statistical Details
While our pairwise bootstrapped p-values are precise, our statistical tests do ignore one potentially important source of imprecision. We assume that each question is independent of the other questions. In a world where forecasting accuracy on questions is quite correlated within a topic, or where most events are correlated, this may result in us overstating how confidently we can reject the equivalence of different models. We hope to explore this question in future work, but we are somewhat reassured by the fact that our gathering of forecasting questions from diverse domains and sources makes it unlikely that correlated questions would change our interpretation of these results in any meaningful way.
Our statistical tests are significantly more precise than the 95% confidence intervals for each model would imply because accuracy on each question is quite correlated across models. To understand this phenomenon, consider a hypothetical world where Model A outperforms Model B by a constant ($\epsilon$) on each question. No matter how close the performance of Models A and B are (how small $\epsilon$ is), and no matter how much variance there is in the accuracy of Model A across questions (which drives the 95% confidence interval surrounding Model A’s accuracy), a pairwise bootstrap would show that Model A is more accurate than Model B. This is because forecasts of Model A and Model B are perfectly correlated.
Appendix H Top 50 ForecastBench Leaderboards
We show the Top 50 leaderboard with human performance evaluated on 200 questions received by humans in Table 17 and the one excluding humans evaluated on the full 1,000 questions in Table 18. Additionally, we show the leaderboard with the inclusion of combination questions in Table 19 , where we generate human forecasts on the combination questions by treating the human forecasts on each component of a combination question as independent (multipling them together to produce the forecast on the relevant combination question). Lastly, we show the leaderboard with combination questions Table 20, where we used the same human-generated forecasts as above but also generated combination forecasts for LLMs from their forecasts on the individual questions, assuming independence of their forecasts on each component, and ignoring the forecast they actually gave on the combination question.
Table 17: Leaderboard with Human Performance: Top 50
% Brier Score $\downarrow$ Model Organization Information provided Prompt Dataset (N=210) Market (N=77) Overall (N=287) Confidence Interval Pairwise p-value comparing to No. 1 Pct. more accurate than No. 1 Superforecaster median forecast ForecastBench 0.135 0.051 0.093 [0.073, 0.112] 0% Public median forecast ForecastBench 0.167 0.048 0.107 [0.09, 0.125] 0.009 30% Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.155 0.067 0.111 [0.09, 0.132] 0.01 33% GPT-4 OpenAI Freeze values Zero shot 0.169 0.056 0.113 [0.093, 0.132] 0.003 37% Claude-3-5-Sonnet-20240620 Anthropic Freeze values Zero shot 0.155 0.074 0.114 [0.09, 0.139] 0.008 36% Claude-3-5-Sonnet-20240620 Anthropic News with freeze values Scratchpad 0.16 0.07 0.115 [0.095, 0.135] 0.004 31% GPT-4o OpenAI Freeze values Scratchpad 0.173 0.059 0.116 [0.098, 0.133] 0.002 30% GPT-4-Turbo-2024-04-09 OpenAI Freeze values Zero shot 0.177 0.055 0.116 [0.097, 0.135] <0.001 37% GPT-4o OpenAI News with freeze values Scratchpad 0.179 0.059 0.119 [0.101, 0.137] <0.001 28% Mistral-Large-Latest Mistral AI Freeze values Scratchpad 0.163 0.081 0.122 [0.107, 0.137] <0.001 24% Mistral-Large-Latest Mistral AI Freeze values Zero shot 0.173 0.074 0.123 [0.102, 0.145] <0.001 28% Claude-3-5-Sonnet-20240620 Anthropic No Scratchpad 0.155 0.093 0.124 [0.103, 0.145] <0.001 29% Llama-3-70b-Chat-Hf Meta Freeze values Zero shot 0.175 0.076 0.125 [0.104, 0.146] <0.001 30% Claude-3-5-Sonnet-20240620 Anthropic News No 0.171 0.082 0.126 [0.108, 0.145] <0.001 28% Claude-3-5-Sonnet-20240620 Anthropic News No 0.167 0.089 0.128 [0.107, 0.148] <0.001 28% Claude-3-Opus-20240229 Anthropic Freeze values Zero shot 0.185 0.071 0.128 [0.106, 0.151] <0.001 33% GPT-4 OpenAI Freeze values Scratchpad 0.188 0.069 0.128 [0.114, 0.143] <0.001 24% Claude-3-5-Sonnet-20240620 Anthropic News Scratchpad 0.16 0.097 0.128 [0.108, 0.149] <0.001 26% GPT-4o OpenAI No Scratchpad 0.173 0.084 0.128 [0.111, 0.146] <0.001 26% Claude-2.1 Anthropic Freeze values Scratchpad 0.206 0.055 0.131 [0.116, 0.145] <0.001 30% Qwen1.5-110B-Chat Qwen Freeze values Zero shot 0.196 0.065 0.131 [0.113, 0.149] <0.001 29% Llama-3-70b-Chat-Hf Meta Freeze values Scratchpad 0.188 0.076 0.132 [0.117, 0.147] <0.001 24% Gemini-1.5-Pro Google Freeze values Zero shot 0.19 0.076 0.133 [0.109, 0.157] <0.001 31% GPT-4o OpenAI Freeze values Zero shot 0.205 0.063 0.134 [0.112, 0.156] <0.001 35% Claude-3-Opus-20240229 Anthropic Freeze values Scratchpad 0.176 0.092 0.134 [0.115, 0.152] <0.001 26% GPT-4o OpenAI News Scratchpad 0.179 0.089 0.134 [0.113, 0.155] <0.001 25% Claude-3-5-Sonnet-20240620 Anthropic No Zero shot 0.155 0.114 0.134 [0.109, 0.16] <0.001 28% Gemini-1.5-Flash Google Freeze values Scratchpad 0.195 0.077 0.136 [0.117, 0.155] <0.001 25% Gemini-1.5-Flash Google Freeze values Zero shot 0.183 0.089 0.136 [0.11, 0.163] <0.001 30% GPT-4-Turbo-2024-04-09 OpenAI No Zero shot 0.177 0.098 0.137 [0.118, 0.156] <0.001 22% Mixtral-8x22B-Instruct-V0.1 Mistral AI Freeze values Zero shot 0.195 0.079 0.137 [0.114, 0.16] <0.001 32% Claude-3-Opus-20240229 Anthropic No Scratchpad 0.176 0.1 0.138 [0.121, 0.155] <0.001 24% GPT-4-Turbo-2024-04-09 OpenAI Freeze values Scratchpad 0.191 0.085 0.138 [0.116, 0.16] <0.001 27% Mixtral-8x22B-Instruct-V0.1 Mistral AI Freeze values Scratchpad 0.19 0.086 0.138 [0.122, 0.154] <0.001 25% Claude-2.1 Anthropic No Scratchpad 0.206 0.07 0.138 [0.121, 0.156] <0.001 28% Gemini-1.5-Pro Google News with freeze values Scratchpad 0.181 0.096 0.138 [0.12, 0.157] <0.001 25% Gemini-1.5-Pro Google Freeze values Scratchpad 0.179 0.099 0.139 [0.121, 0.157] <0.001 26% GPT-4 OpenAI No Scratchpad 0.188 0.096 0.142 [0.128, 0.157] <0.001 20% Mistral-Large-Latest Mistral AI No Scratchpad 0.163 0.123 0.143 [0.126, 0.16] <0.001 24% GPT-4-Turbo-2024-04-09 OpenAI No Scratchpad 0.191 0.095 0.143 [0.126, 0.16] <0.001 23% Imputed Forecaster ForecastBench 0.25 0.037 0.143 [0.129, 0.158] <0.001 34% Claude-2.1 Anthropic Freeze values Zero shot 0.22 0.067 0.144 [0.124, 0.163] <0.001 31% Gemini-1.5-Pro Google No Scratchpad 0.179 0.109 0.144 [0.125, 0.163] <0.001 24% GPT-4 OpenAI No Zero shot 0.169 0.119 0.144 [0.127, 0.162] <0.001 23% Gemini-1.5-Pro Google News Scratchpad 0.181 0.112 0.146 [0.127, 0.166] <0.001 26% GPT-4-Turbo-2024-04-09 OpenAI News with freeze values Scratchpad 0.202 0.09 0.146 [0.124, 0.169] <0.001 31% Llama-3-70b-Chat-Hf Meta No Zero shot 0.175 0.12 0.148 [0.128, 0.168] <0.001 24% Claude-3-Opus-20240229 Anthropic No Zero shot 0.185 0.112 0.149 [0.125, 0.173] <0.001 25% Qwen1.5-110B-Chat Qwen News with freeze values Scratchpad 0.196 0.103 0.149 [0.132, 0.167] <0.001 25% Mixtral-8x7B-Instruct-V0.1 Mistral AI Freeze values Zero shot 0.209 0.09 0.15 [0.12, 0.179] <0.001 36%
Notes:
% reduces space between this line and the next one Pairwise p-value comparing to No. 1 (bootstrapped): The p-value calculated by bootstrapping the differences in overall score between each model and the best forecaster (the group with rank 1) under the null hypothesis that there’s no difference.
% reduces space between this line and the next one Pct. more accurate than No. 1: The percent of questions where this forecaster had a better overall score than the best forecaster (with rank 1).
Table 18: Leaderboard excluding Human Performance: Top 50
% Brier Score $\downarrow$ Model Organization Information provided Prompt Dataset (N=210) Market (N=77) Overall (N=287) Confidence Interval Pairwise p-value comparing to No. 1 Pct. more accurate than No. 1 GPT-4-Turbo-2024-04-09 OpenAI Freeze values Scratchpad 0.175 0.051 0.113 [0.107, 0.118] 0% Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.179 0.047 0.113 [0.108, 0.119] 0.442 53% GPT-4o OpenAI Freeze values Scratchpad 0.201 0.039 0.12 [0.115, 0.125] <0.001 45% Gemini-1.5-Pro Google Freeze values Scratchpad 0.168 0.075 0.121 [0.116, 0.127] <0.001 42% GPT-4o OpenAI News with freeze values Scratchpad 0.196 0.05 0.123 [0.117, 0.128] <0.001 41% Claude-3-Opus-20240229 Anthropic Freeze values Zero shot 0.192 0.061 0.127 [0.12, 0.133] <0.001 44% Gemini-1.5-Pro Google News with freeze values Scratchpad 0.172 0.082 0.127 [0.121, 0.132] <0.001 41% GPT-4-Turbo-2024-04-09 OpenAI Freeze values Zero shot 0.199 0.055 0.127 [0.12, 0.133] <0.001 47% Qwen1.5-110B-Chat Qwen Freeze values Scratchpad 0.176 0.078 0.127 [0.122, 0.133] <0.001 32% GPT-4-Turbo-2024-04-09 OpenAI No Scratchpad 0.175 0.083 0.129 [0.124, 0.134] <0.001 7% Claude-3-5-Sonnet-20240620 Anthropic News with freeze values Scratchpad 0.197 0.067 0.132 [0.127, 0.138] <0.001 43% Gemini-1.5-Pro Google No Scratchpad 0.168 0.097 0.133 [0.127, 0.138] <0.001 39% GPT-4 OpenAI Freeze values Scratchpad 0.198 0.067 0.133 [0.127, 0.138] <0.001 33% Claude-3-5-Sonnet-20240620 Anthropic No Scratchpad 0.179 0.087 0.133 [0.127, 0.139] <0.001 46% Llama-3-70b-Chat-Hf Meta Freeze values Zero shot 0.199 0.069 0.134 [0.128, 0.14] <0.001 38% Gemini-1.5-Pro Google News Scratchpad 0.172 0.097 0.134 [0.128, 0.14] <0.001 39% Gemini-1.5-Pro Google Freeze values Zero shot 0.219 0.052 0.135 [0.129, 0.141] <0.001 44% Claude-3-5-Sonnet-20240620 Anthropic Freeze values Zero shot 0.208 0.063 0.136 [0.129, 0.142] <0.001 51% Imputed Forecaster ForecastBench 0.25 0.021 0.136 [0.133, 0.138] <0.001 43% LLM Crowd ForecastBench News 0.234 0.039 0.137 [0.134, 0.14] <0.001 37% GPT-4-Turbo-2024-04-09 OpenAI News with freeze values Scratchpad 0.213 0.061 0.137 [0.132, 0.142] <0.001 29% Gemini-1.5-Flash Google Freeze values Scratchpad 0.187 0.088 0.138 [0.131, 0.144] <0.001 40% GPT-4o OpenAI News Scratchpad 0.196 0.082 0.139 [0.133, 0.145] <0.001 34% Qwen1.5-110B-Chat Qwen No Scratchpad 0.176 0.102 0.139 [0.134, 0.144] <0.001 30% LLM Crowd ForecastBench News 0.236 0.042 0.139 [0.136, 0.142] <0.001 36% LLM Crowd ForecastBench News 0.237 0.042 0.139 [0.136, 0.142] <0.001 36% GPT-4 OpenAI Freeze values Zero shot 0.222 0.057 0.14 [0.134, 0.145] <0.001 38% Mistral-Large-Latest Mistral AI Freeze values Scratchpad 0.197 0.082 0.14 [0.135, 0.145] <0.001 29% GPT-4o OpenAI No Scratchpad 0.201 0.081 0.141 [0.136, 0.147] <0.001 37% Llama-3-70b-Chat-Hf Meta Freeze values Scratchpad 0.208 0.075 0.142 [0.137, 0.147] <0.001 29% GPT-4-Turbo-2024-04-09 OpenAI No Zero shot 0.199 0.086 0.142 [0.136, 0.148] <0.001 38% Gemini-1.5-Pro Google No Zero shot 0.219 0.069 0.144 [0.138, 0.15] <0.001 42% Claude-3-5-Sonnet-20240620 Anthropic News Scratchpad 0.197 0.091 0.144 [0.139, 0.15] <0.001 40% GPT-4 OpenAI No Scratchpad 0.198 0.091 0.145 [0.14, 0.149] <0.001 29% Claude-3-Opus-20240229 Anthropic Freeze values Scratchpad 0.21 0.08 0.145 [0.139, 0.151] <0.001 36% Mistral-Large-Latest Mistral AI Freeze values Zero shot 0.201 0.091 0.146 [0.139, 0.153] <0.001 35% Gemini-1.5-Pro Google News No 0.193 0.099 0.146 [0.14, 0.152] <0.001 34% Gemini-1.5-Flash Google No Scratchpad 0.187 0.106 0.147 [0.141, 0.153] <0.001 37% Gemini-1.5-Flash Google Freeze values Zero shot 0.225 0.068 0.147 [0.14, 0.154] <0.001 45% Llama-3-70b-Chat-Hf Meta No Zero shot 0.199 0.095 0.147 [0.141, 0.152] <0.001 34% Claude-3-5-Sonnet-20240620 Anthropic News No 0.199 0.096 0.148 [0.142, 0.153] <0.001 38% Claude-3-Opus-20240229 Anthropic No Zero shot 0.192 0.104 0.148 [0.141, 0.155] <0.001 37% Qwen1.5-110B-Chat Qwen News with freeze values Scratchpad 0.207 0.09 0.149 [0.143, 0.154] <0.001 27% GPT-4-Turbo-2024-04-09 OpenAI News Scratchpad 0.213 0.085 0.149 [0.144, 0.154] <0.001 24% GPT-4-Turbo-2024-04-09 OpenAI News No 0.212 0.086 0.149 [0.144, 0.154] <0.001 27% Mixtral-8x22B-Instruct-V0.1 Mistral AI Freeze values Scratchpad 0.214 0.091 0.152 [0.146, 0.158] <0.001 28% Claude-2.1 Anthropic No Scratchpad 0.232 0.072 0.152 [0.148, 0.157] <0.001 36% GPT-4o OpenAI News No 0.212 0.093 0.152 [0.147, 0.158] <0.001 28% Claude-3-5-Sonnet-20240620 Anthropic No Zero shot 0.208 0.1 0.154 [0.147, 0.161] <0.001 43% Claude-3-5-Sonnet-20240620 Anthropic News No 0.213 0.096 0.154 [0.149, 0.16] <0.001 34%
Notes:
% reduces space between this line and the next one Pairwise p-value comparing to No. 1 (bootstrapped): The p-value calculated by bootstrapping the differences in overall score between each model and the best forecaster (the group with rank 1) under the null hypothesis that there’s no difference.
% reduces space between this line and the next one Pct. more accurate than No. 1: The percent of questions where this forecaster had a better overall score than the best forecaster (with rank 1).
Table 19: Leaderboard with Performance om Combination Questions: Top 50
% Brier Score $\downarrow$ Model Organization Information provided Prompt Dataset (N=210) Market (N=77) Overall (N=287) Confidence Interval Pairwise p-value comparing to No. 1 Pct. more accurate than No. 1 Superforecaster median forecast ForecastBench 0.097 0.04 0.068 [0.06, 0.077] 0% Public median forecast ForecastBench 0.127 0.039 0.083 [0.075, 0.091] <0.001 28% Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.168 0.069 0.119 [0.107, 0.13] <0.001 27% GPT-4o OpenAI Freeze values Scratchpad 0.191 0.048 0.119 [0.11, 0.129] <0.001 28% GPT-4o OpenAI News with freeze values Scratchpad 0.179 0.06 0.119 [0.11, 0.129] <0.001 27% GPT-4-Turbo-2024-04-09 OpenAI Freeze values Scratchpad 0.171 0.068 0.12 [0.109, 0.131] <0.001 27% Gemini-1.5-Pro Google Freeze values Scratchpad 0.16 0.092 0.126 [0.116, 0.137] <0.001 24% GPT-4-Turbo-2024-04-09 OpenAI Freeze values Zero shot 0.193 0.064 0.128 [0.117, 0.139] <0.001 30% Claude-3-5-Sonnet-20240620 Anthropic News with freeze values Scratchpad 0.173 0.088 0.131 [0.119, 0.142] <0.001 23% Claude-3-Opus-20240229 Anthropic Freeze values Zero shot 0.183 0.083 0.133 [0.12, 0.145] <0.001 27% Gemini-1.5-Pro Google News with freeze values Scratchpad 0.161 0.104 0.133 [0.122, 0.144] <0.001 22% Qwen1.5-110B-Chat Qwen Freeze values Scratchpad 0.173 0.093 0.133 [0.123, 0.143] <0.001 18% GPT-4-Turbo-2024-04-09 OpenAI No Scratchpad 0.171 0.096 0.134 [0.124, 0.143] <0.001 19% Claude-3-5-Sonnet-20240620 Anthropic Freeze values Zero shot 0.189 0.079 0.134 [0.122, 0.147] <0.001 26% GPT-4o OpenAI News Scratchpad 0.179 0.093 0.136 [0.125, 0.147] <0.001 21% Gemini-1.5-Pro Google No Scratchpad 0.16 0.112 0.136 [0.126, 0.147] <0.001 21% GPT-4 OpenAI Freeze values Scratchpad 0.192 0.08 0.136 [0.126, 0.146] <0.001 21% Claude-3-5-Sonnet-20240620 Anthropic No Scratchpad 0.168 0.105 0.137 [0.125, 0.148] <0.001 21% Llama-3-70b-Chat-Hf Meta Freeze values Zero shot 0.188 0.089 0.138 [0.127, 0.15] <0.001 23% GPT-4 OpenAI Freeze values Zero shot 0.213 0.064 0.139 [0.128, 0.149] <0.001 27% Gemini-1.5-Pro Google Freeze values Zero shot 0.208 0.069 0.139 [0.127, 0.15] <0.001 26% GPT-4o OpenAI No Scratchpad 0.191 0.086 0.139 [0.128, 0.149] <0.001 20% Gemini-1.5-Pro Google News Scratchpad 0.161 0.116 0.139 [0.127, 0.15] <0.001 21% Imputed Forecaster ForecastBench 0.25 0.029 0.14 [0.134, 0.145] <0.001 29% LLM Crowd ForecastBench News 0.232 0.048 0.14 [0.134, 0.146] <0.001 24% Gemini-1.5-Flash Google Freeze values Scratchpad 0.175 0.107 0.141 [0.129, 0.153] <0.001 22% Claude-3-5-Sonnet-20240620 Anthropic News Scratchpad 0.173 0.11 0.142 [0.13, 0.153] <0.001 20% LLM Crowd ForecastBench News 0.234 0.051 0.142 [0.136, 0.148] <0.001 24% LLM Crowd ForecastBench News 0.236 0.05 0.143 [0.137, 0.149] <0.001 24% Llama-3-70b-Chat-Hf Meta Freeze values Scratchpad 0.196 0.089 0.143 [0.134, 0.152] <0.001 18% Claude-3-Opus-20240229 Anthropic Freeze values Scratchpad 0.192 0.095 0.143 [0.133, 0.154] <0.001 19% Mistral-Large-Latest Mistral AI Freeze values Scratchpad 0.186 0.102 0.144 [0.134, 0.153] <0.001 18% Gemini-1.5-Flash Google No Scratchpad 0.175 0.114 0.144 [0.134, 0.155] <0.001 17% Claude-3-5-Sonnet-20240620 Anthropic News No 0.178 0.111 0.145 [0.133, 0.156] <0.001 19% GPT-4-Turbo-2024-04-09 OpenAI News with freeze values Scratchpad 0.21 0.08 0.145 [0.134, 0.156] <0.001 24% Qwen1.5-110B-Chat Qwen No Scratchpad 0.173 0.119 0.146 [0.137, 0.155] <0.001 16% GPT-4-Turbo-2024-04-09 OpenAI No Zero shot 0.193 0.1 0.146 [0.134, 0.158] <0.001 21% Gemini-1.5-Pro Google No Zero shot 0.208 0.087 0.148 [0.136, 0.16] <0.001 21% GPT-4 OpenAI No Scratchpad 0.192 0.107 0.149 [0.141, 0.158] <0.001 16% Claude-2.1 Anthropic No Scratchpad 0.224 0.076 0.15 [0.141, 0.159] <0.001 24% Gemini-1.5-Flash Google Freeze values Zero shot 0.216 0.084 0.15 [0.136, 0.163] <0.001 27% Mistral-Large-Latest Mistral AI Freeze values Zero shot 0.197 0.104 0.15 [0.138, 0.163] <0.001 22% Gemini-1.5-Pro Google News No 0.181 0.122 0.151 [0.14, 0.163] <0.001 20% Claude-3-Opus-20240229 Anthropic No Zero shot 0.183 0.122 0.152 [0.139, 0.165] <0.001 22% GPT-4-Turbo-2024-04-09 OpenAI News Scratchpad 0.21 0.097 0.153 [0.143, 0.164] <0.001 19% Qwen1.5-110B-Chat Qwen News with freeze values Scratchpad 0.199 0.109 0.154 [0.144, 0.165] <0.001 18% Llama-3-70b-Chat-Hf Meta No Zero shot 0.188 0.121 0.154 [0.144, 0.165] <0.001 17% Claude-2.1 Anthropic Freeze values Scratchpad 0.224 0.089 0.156 [0.146, 0.167] <0.001 23% GPT-4-Turbo-2024-04-09 OpenAI News No 0.206 0.108 0.157 [0.147, 0.168] <0.001 19% Mistral-Large-Latest Mistral AI No Scratchpad 0.186 0.131 0.158 [0.149, 0.168] <0.001 16%
Notes:
% reduces space between this line and the next one Pairwise p-value comparing to No. 1 (bootstrapped): The p-value calculated by bootstrapping the differences in overall score between each model and the best forecaster (the group with rank 1) under the null hypothesis that there’s no difference.
% reduces space between this line and the next one Pct. more accurate than No. 1: The percent of questions where this forecaster had a better overall score than the best forecaster (with rank 1).
Table 20: Leaderboard with Performance on Combination Questions (Independence): Top 50
Brier Score $\downarrow$
Model Organization
Information
provided
Prompt
Dataset
(N=870)
Market
(N=301)
Overall
(N=1,171)
Confidence
Interval
Pairwise
p-value
comparing
to No. 1
Pct. more
accurate
than No. 1
Superforecaster median forecast ForecastBench 0.097 0.04 0.068 [0.06, 0.077] 0%
Public median forecast ForecastBench 0.127 0.039 0.083 [0.075, 0.091] $<$0.001 28%
GPT-4 OpenAI Freeze values Zero shot 0.131 0.041 0.086 [0.077, 0.095] $<$0.001 34%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Zero shot 0.135 0.04 0.088 [0.079, 0.097] $<$0.001 33%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Zero shot 0.114 0.062 0.088 [0.076, 0.1] $<$0.001 34%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.118 0.064 0.091 [0.08, 0.102] $<$0.001 28%
GPT-4o OpenAI Freeze values Scratchpad 0.134 0.052 0.093 [0.084, 0.102] $<$0.001 28%
Claude-3-5-Sonnet-20240620 Anthropic News with freeze values Scratchpad 0.13 0.06 0.095 [0.085, 0.105] $<$0.001 28%
GPT-4o OpenAI News with freeze values Scratchpad 0.142 0.049 0.096 [0.087, 0.104] $<$0.001 27%
Mistral-Large-Latest Mistral AI Freeze values Zero shot 0.134 0.062 0.098 [0.088, 0.108] $<$0.001 25%
Mistral-Large-Latest Mistral AI Freeze values Scratchpad 0.13 0.068 0.099 [0.091, 0.107] $<$0.001 24%
Claude-3-Opus-20240229 Anthropic Freeze values Zero shot 0.146 0.057 0.102 [0.091, 0.112] $<$0.001 30%
Claude-3-5-Sonnet-20240620 Anthropic No Scratchpad 0.118 0.086 0.102 [0.09, 0.113] $<$0.001 26%
Claude-2.1 Anthropic Freeze values Scratchpad 0.163 0.042 0.102 [0.095, 0.11] $<$0.001 29%
Llama-3-70b-Chat-Hf Meta Freeze values Zero shot 0.139 0.067 0.103 [0.093, 0.114] $<$0.001 26%
GPT-4o OpenAI No Scratchpad 0.134 0.073 0.104 [0.094, 0.113] $<$0.001 24%
Claude-3-5-Sonnet-20240620 Anthropic News No 0.135 0.074 0.104 [0.094, 0.114] $<$0.001 25%
Qwen1.5-110B-Chat Qwen Freeze values Zero shot 0.153 0.056 0.105 [0.096, 0.114] $<$0.001 26%
Claude-3-Opus-20240229 Anthropic Freeze values Scratchpad 0.135 0.076 0.105 [0.096, 0.115] $<$0.001 24%
GPT-4 OpenAI Freeze values Scratchpad 0.151 0.061 0.106 [0.097, 0.114] $<$0.001 24%
Claude-3-5-Sonnet-20240620 Anthropic News Scratchpad 0.13 0.083 0.107 [0.096, 0.118] $<$0.001 24%
Claude-2.1 Anthropic No Scratchpad 0.163 0.052 0.107 [0.099, 0.116] $<$0.001 27%
Claude-3-5-Sonnet-20240620 Anthropic No Zero shot 0.114 0.101 0.107 [0.094, 0.12] $<$0.001 26%
Llama-3-70b-Chat-Hf Meta Freeze values Scratchpad 0.151 0.064 0.108 [0.099, 0.116] $<$0.001 23%
Mixtral-8x22B-Instruct-V0.1 Mistral AI Freeze values Zero shot 0.154 0.066 0.11 [0.099, 0.121] $<$0.001 29%
Gemini-1.5-Flash Google Freeze values Scratchpad 0.151 0.069 0.11 [0.1, 0.12] $<$0.001 25%
Claude-3-5-Sonnet-20240620 Anthropic News No 0.136 0.085 0.11 [0.099, 0.121] $<$0.001 24%
GPT-4o OpenAI Freeze values Zero shot 0.166 0.055 0.11 [0.1, 0.121] $<$0.001 30%
GPT-4o OpenAI News Scratchpad 0.142 0.08 0.111 [0.1, 0.122] $<$0.001 23%
Gemini-1.5-Pro Google Freeze values Zero shot 0.154 0.068 0.111 [0.099, 0.123] $<$0.001 28%
Gemini-1.5-Pro Google Freeze values Scratchpad 0.139 0.084 0.111 [0.102, 0.121] $<$0.001 23%
Claude-3-Opus-20240229 Anthropic No Scratchpad 0.135 0.088 0.111 [0.102, 0.121] $<$0.001 23%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Scratchpad 0.152 0.072 0.112 [0.101, 0.122] $<$0.001 26%
GPT-4-Turbo-2024-04-09 OpenAI No Zero shot 0.135 0.089 0.112 [0.102, 0.123] $<$0.001 22%
GPT-4 OpenAI No Zero shot 0.131 0.095 0.113 [0.103, 0.122] $<$0.001 24%
Gemini-1.5-Pro Google News with freeze values Scratchpad 0.143 0.084 0.113 [0.103, 0.123] $<$0.001 22%
Mistral-Large-Latest Mistral AI No Scratchpad 0.13 0.097 0.113 [0.104, 0.123] $<$0.001 23%
Mixtral-8x22B-Instruct-V0.1 Mistral AI Freeze values Scratchpad 0.15 0.077 0.114 [0.105, 0.123] $<$0.001 23%
Gemini-1.5-Flash Google Freeze values Zero shot 0.15 0.078 0.114 [0.101, 0.127] $<$0.001 27%
Gemini-1.5-Pro Google No Scratchpad 0.139 0.09 0.115 [0.105, 0.124] $<$0.001 22%
GPT-4 OpenAI No Scratchpad 0.151 0.079 0.115 [0.106, 0.123] $<$0.001 21%
Imputed Forecaster ForecastBench 0.203 0.029 0.116 [0.107, 0.124] $<$0.001 32%
Gemini-1.5-Pro Google News Scratchpad 0.143 0.091 0.117 [0.107, 0.127] $<$0.001 23%
GPT-4-Turbo-2024-04-09 OpenAI No Scratchpad 0.152 0.082 0.117 [0.108, 0.127] $<$0.001 21%
Claude-2.1 Anthropic Freeze values Zero shot 0.179 0.055 0.117 [0.107, 0.127] $<$0.001 29%
GPT-4-Turbo-2024-04-09 OpenAI News with freeze values Scratchpad 0.161 0.075 0.118 [0.108, 0.129] $<$0.001 26%
Claude-3-Opus-20240229 Anthropic No Zero shot 0.146 0.093 0.119 [0.108, 0.131] $<$0.001 24%
Claude-3-Opus-20240229 Anthropic News No 0.134 0.106 0.12 [0.11, 0.131] $<$0.001 23%
Llama-3-70b-Chat-Hf Meta No Zero shot 0.139 0.102 0.12 [0.11, 0.131] $<$0.001 23%
Gemini-1.5-Flash Google No Scratchpad 0.151 0.09 0.12 [0.11, 0.13] $<$0.001 22%
Notes:
Pairwise p-value comparing to No. 1 (bootstrapped): The p-value calculated by bootstrapping the differences in overall score between each model and the best forecaster (the group with rank 1) under the null hypothesis that there’s no difference.
Pct. more accurate than No. 1: The percent of questions where this forecaster had a better overall score than the best forecaster (with rank 1).